OpenStudy (anonymous):

Differentiate f(x)=ln(x2+6).

OpenStudy (anonymous):

do you know the chain rule?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so look at it like this: \[d(ln(x^{2}+6))/dx\\\\=d (ln(something))/dx\\ =\frac{1}{something}*\frac{d(something)}{dx}\]

OpenStudy (precal):